"N=2^{12}=4096" ways

How to review the solution in polya strategy


1
Expert's answer
2021-10-27T06:30:15-0400

In how many ways can you answer a 12-question true-false test if you answer each question with either a “true” or a “false”?

Solution

Step 1: Understand the problem.

Step 2: Devise a plan (translate).

Step 3: Carry out the plan (solve).

Step 4: Look back (check and interpret).


Step 1:Since there are 12 questions, and each question can be answered in 2 ways then:

1st question can be answered in 2 ways.

2nd question can be answered in 2 ways.


Step 2: The two questions can be answered in "2\times2=2^2" ways.

The three questions can be answered in "2\times2\times 2=2^3" ways. and so on.

Step 3: So we have:


"N=2^{12}=4096"

Step 4: There are 4096 ways to answer a 12-question true-false test if you answer each question with either a “true” or a “false”.

Therefore we have to prepare before the test to obtain the good score.